Solve by completing the square and quadratic formula:
1.) x^2-6x+8=0
2.)t^2=4-3t

Answer by mananth(16946)


x^2-6x+8=0
x^2-6x = -8
x^2-6x+9= +9-8
(x-3)^2= 1
take the square root
(x-3) = +/-1
x=4 OR 2
t^2=4-3t
t^2+3t =4
Third term = ((1/2) the co-efficient of second term)^2
t^2+3t + 9/4 = 4+9/4
(t+ 3/2)^2= 25/4
t+ 3/2 = +/- 5/2
t = -3/2 + 4/2 OR -3/2 -5/2
t= 1/2 OR -4